OpenStudy (anonymous):

f(x,y) = x+2y. Evaluate scalar line integral over given path x. x(t)=(2-3t,4t-1), 0

OpenStudy (anonymous):

wrt dx or dy or ds?

OpenStudy (anonymous):

ds

OpenStudy (anonymous):

the answer is 50, but im getting 54

OpenStudy (anonymous):

(2-3t + 8t - 2)ds 5tds ds = sqrt(25)

OpenStudy (anonymous):

so u get 25tdt from t=0 to 2 get it?

OpenStudy (anonymous):

sorry, you lost me at previous post

OpenStudy (anonymous):

ohk wts the formula for ds?

OpenStudy (anonymous):

my teacher uses the equation \[\int\limits_{0}^{2} f(x(t))*(x'(t)) ds\]

OpenStudy (anonymous):

so for this problem i got \[\int\limits_{0}^{2} (2-3t,4t-1) * (-3,4) dt\]

OpenStudy (anonymous):

er should be (2-3t,8t-2)

OpenStudy (anonymous):

is the answer 50??

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

then look ure supposed to evaluate f(x,y)ds arent u?

OpenStudy (anonymous):

fox????????

OpenStudy (anonymous):

convert f(x,y) to f(t) by replacing x and y by their parametric definitions ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so f(x,y) reduces to what?

OpenStudy (anonymous):

(2-3t,8t-2) ?

OpenStudy (anonymous):

x+2y = 5t

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

isnt it??

OpenStudy (anonymous):

so now for ds remember \[ds = \sqrt{ (dx/dt)^2 + (dy/dt)^2} dt\]

OpenStudy (anonymous):

so, sqrt(3^2+4^2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which =5

OpenStudy (anonymous):

so finally u have \[\int\limits_{0}^{2}(5t)(5dt)\]

OpenStudy (anonymous):

now its easy isnt it??

OpenStudy (anonymous):

yea. thanks

OpenStudy (anonymous):

welcome